# Solving dynamic portfolio selection problems via score-based diffusion models

### v1

@@ Abstract
In this paper, we tackle the dynamic mean-variance portfolio selection problem in a model-free manner, based on (generative) diffusion models. We propose using data sampled from the real model $\mathbb{P}$ (which is unknown) with limited size to train a generative model $\mathbb{Q}$ (from which we can easily and adequately sample). With adaptive training and sampling methods that are tailor-made for time series data, we obtain quantification bounds between $\mathbb{P}$ and $\mathbb{Q}$ in terms of the adapted Wasserstein metric $\mathcal{A}\mathcal{W}_{2}$. Importantly, the proposed adapted sampling method also facilitates conditional sampling. In the second part of this paper, we provide the stability of the mean-variance portfolio optimization problems in $\mathcal{A}\mathcal{W}_{2}$. Then, combined with the error bounds and the stability result, we propose a policy gradient algorithm based on the generative environment, in which our innovative adapted sampling method provides approximate scenario generators. We illustrate the performance of our algorithm on both simulated and real data. For real data, the algorithm based on the generative environment produces portfolios that beat several important baselines, including the Markowitz portfolio, the equal weight (naive) portfolio, and S&P 500.
@@ 1 Introduction
Portfolio selection, a crucial and fundamental task in financial engineering, has been a subject of both theoretical and practical interest since the classical work Markowitz (1952). With the ability to account for intertemporal hedging and reblancing in a nonstationary market, formulating portfolio selection as dynamic multiperiod problems has become increasingly appealing; see Mossin (1968), Merton (1971) for early works. Researchers have explored this problem under various specifications, such as continuous-time dynamic mean-variance problems, where return models are prescribed by some stochastic differential equations Zhou and Li (2000), or discrete-time problems with independent intertemporal returns Li and Ng (2000). However, due to the incompleteness of the information, estimation errors, or simply the infeasibility of fully modeling market uncertainty, we lack access to the real model, as opposed to model-based approaches. Instead, we are restricted to a limited set of financial data samples. Consequently, investigating portfolio selection problems in a model-free manner has been a long-standing topic in financial engineering; see discussions on ‘universal portfolio” Cover (1991), stochastic portfolio theory Karatzas and Ruf (2017), Cuchiero et al. (2019), and the application of machine learning Ban et al. (2018), Gu et al. (2020). However, dynamic portfolio selection problems are rather involved, even with model-based approaches, due to their inherent intertemporal structure Campbell and Viceira (1999), high dimensionality Brandt et al. (2005) and time inconsistency Basak and Chabakauri (2010), Björk et al. (2014). Therefore, model-free dynamic portfolio selection problems are worth exploring.
@@
To address this challenge, we propose to leverage score-based diffusion models. These models can generate additional samples that exhibit distributional properties similar to the original data. Specifically, with an “unknown” model $s^{1:T}\sim\mathbb{P}$, we propose a data-driven approach to train a model $\mathbb{Q}$ that closely resembles $\mathbb{P}$. We utilize the model $\mathbb{Q}$ as an oracle in the subsequent policy gradient algorithm: we sample a large number of paths $\tilde{s}^{1:T}\sim\mathbb{Q}$ and train a reinforcement learning (RL) agent using these simulated paths. Theoretically, we show that our method achieves near-optimality in terms of Markowitz’s mean-variance criterion, thanks to two key contributions of the present paper: (1) a quantification bound between $\mathbb{P}$ and $\mathbb{Q}$, and (2) the model stability. From an engineering perspective, the approximated model $\mathbb{Q}$ consists of a score network $s_{\theta}$, which is the core of the score-matching technique used in generative models, and a Recurrent Neural Network (RNN) encoder $R_{\theta}$, which encodes the current state of the market. These two networks are trained simultaneously using historical price data. The reinforcement learning agent is trained with the generated data from $\mathbb{Q}$, which are produced by evaluating $s_{\theta}$ and $R_{\theta}$ in an adaptive manner. Finally, the agent, fed with the encoded (and updated) market state, outputs portfolio selection actions whose performance is evaluated through real-world experiments. For a summary of the theoretical results, see Figure 1, and for the workflow when implementing the proposed solution for model-free dynamic portfolio selection, refer to Figure 2.
Y1.0e0]ptMV problemunder $\mathbb{P}$MV problemunder $\mathbb{Q}$Quadratic hedgingunder $\mathbb{Q}$Financial Datafrom $\mathbb{P}$Pretrained $\mathbb{Q}$Policy gradientalgorithm(Algorithm 4) Model Stability (Corollary 3.9) Duality (3.5) Adaptive sampling (Section 2) Quantification error bounds in $\mathcal{A}\mathcal{W}_{2}$ (Section 2) Model-based approach (omniscient solution)
@@ 1.1 Background, main results and contributions
The score matching technique, which has been a powerful tool in statistics since the fundamental work Hyvärinen (2005), has also been incorporated into generative models in various works, including Sohl-Dickstein et al. (2015), Song and Ermon (2019), and others. Specifically, Song et al. (2021b) unifies the concept of score-based generative models and diffusion probabilistic models (Ho et al. (2020a)) into a single framework, known as score-based diffusion models, which we adopt in this work.
@@
where $p(\tau,\cdot)$ denotes the density of $X_{\tau}$. Because $p_{\rm data}$ is unknown, so is $p(\tau,\cdot)$. Under the score‑matching theory, an accurate estimate of the score $\nabla\log p$ guarantees that, for sufficiently large $\mathcal{T}$, the law of $Y_{\mathcal{T}}$ approximates $p_{\rm data}$ under various metrics; see Tang and Zhao (2024b) for a comprehensive survey and Gao et al. (2025) for recent Wasserstein bounds. Score-based diffusion models have found numerous applications across a wide range of areas, including image and video generation (Rombach et al. (2022), Melnik et al. (2024)), molecule simulation Hoogeboom et al. (2022) and health Abramson et al. (2024).
@@
Nevertheless, because we are interested in dynamic decision-making problems, the data take the form of time series: $X=(X^{1},X^{2},\cdots,X^{T})=X^{1:T}$. For instance, $T=12$ and each $X^{t}$ is the price data in a monthly frequency, then $X$ is a one-year price path. Although score-based diffusion models have been applied to time-series prediction (see Lin et al. (2024) for a survey), to the best of our knowledge, there is still no theoretical result that establishes foundations for time-series data generation. Treating the state space naively as the enlarged Euclidean space $\mathbb{R}^{dT}$ is unsatisfactory: it overlooks the intertemporal structure and the non-anticipation requirements inherent in dynamic problems. Indeed, existing results show that, for dynamic settings, value functions are unstable under standard probabilistic metrics such as Wasserstein; see, e.g., Backhoff-Veraguas et al. (2020). This observation motivates the use of the adapted Wasserstein metric, $\mathcal{A}\mathcal{W}_{2}$, which is better suited to time-series data and dynamic tasks; see Backhoff-Veraguas et al. (2020) for definitions. Furthermore, a dynamic control problem requires sampling not only from the joint distribution of $\mathbb{P}$ but also from the conditional distribution $\mathbb{P}(\cdot|X^{1:t}=x^{1:t})$ given observations up to time $t$. Such conditional sampling is indispensable for our subsequent policy-gradient algorithm (i.e., for training the RL agent). Traditional score-based diffusion models, however, do not support this functionality. Consequently, we must develop new training and sampling procedures that (i) admit theoretical score-matching error bounds under the adapted Wasserstein metric and (ii) enable conditional sampling for any observation input.
@@
Our first set of results meets these two requirements. For notational convenience we fix $\beta\equiv 2$ in (1.1) and, for each $t=1,2,\dots,T-1$, consider the following conditional VP-SDE:
$\begin{cases}&\mathrm{d}X^{t+1}_{\tau}=-X^{t+1}_{\tau},\mathrm{d}\tau+\sqrt{2}\mathrm{d}B^{t+1}_{\tau},\\ &X^{t+1}_{0}\sim\mathbb{P}_{x^{1:t}},\end{cases},$ (1.3)
@@
where $\mathbb{P}_{x^{1:t}}$ denotes the law of $X^{t+1}$ conditional on $X^{1:t}=x^{1:t}$. This formulation allows us to define the conditional score function $s^{t+1}(\tau,x^{1:t},x)=\nabla_{x}\log p_{t+1}(\tau,x^{1:t},x)$, where $p_{t+1}(\tau,x^{1:t},\cdot)$ is the density of $X^{t+1}_{\tau}$ in (1.3). Assume we have an approximator $\hat{s}^{t+1}$ that is close to $s$ in a suitable $L^{2}$ sense. We generate a new trajectory $y^{1:T}$ adaptively: to obtain $y^{1}$ we start from pure noise and run the reverse VP-SDE (1.2), setting $y^{1}=Y_{\mathcal{T}}$; to obtain $y^{t+1}$ we use the previously generated path $y^{1;t}$ as conditional input and run (1.2) with the score replaced by $\hat{s}^{t+1}(\cdot,y^{1:t},\cdot)$; see Algorithm 1. Under appropriate technical conditions we obtain the following bound:
@@
Denote by $\mathbb{Q}^{\mathcal{T}}$ the distribution of $y^{1:T}$. Then $\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq E(\mathcal{T},\varepsilon_{\rm score})$, where $\varepsilon_{\rm score}$ is the score-matching error between $\hat{s}$ and $s$. Moreover, $E(\mathcal{T},\varepsilon)$ has an explicit form and can be made arbitrarily small by taking $\mathcal{T}$ sufficiently large and $\varepsilon_{\rm score}$ sufficiently small.
@@
While we can now generate new samples from $\mathbb{Q}$, several steps remain before we can fully address the dynamic portfolio‑selection problem. First, because we now work under the surrogate model $\mathbb{Q}$, we must establish model stability: even if $\mathbb{P}$ and $\mathbb{Q}$ are close and we can solve the mean–variance problem under $\mathbb{Q}$, does the optimizer obtained from $\mathbb{Q}$ remain near‑optimal for the original model $\mathbb{P}$? We answer this question in the affirmative:
@@
    We extend the score‑based diffusion‑model framework to time‑series data. Using a novel adaptive training and sampling scheme, we establish theoretical error bounds under the adapted Wasserstein metric. The framework also supports conditional sampling, which is crucial for solving model‑free portfolio‑selection problems and for constructing high‑performance portfolio strategies;
@@ 1.2 Related works
The practice of simulating artificial price paths to guide investment decisions predates modern portfolio theory. Early random walk studies of convertible warrant pricing by Sprenkle (1961) and Samuelson (1965) already relied on Monte Carlo draws, while Hertz (1964) popularized the technique in corporate finance capital budgeting. During the 1970s–1990s, more sophisticated parametric dynamics, jumping diffusion Merton (1976), stochastic volatility Heston (1993), and SV with jumps Bates (1996) - made large-scale pathwise risk analysis computationally tractable.
@@
As asset universes became high-dimensional and tail dependencies more intricate, researchers began exploring fully data-driven generators. Early adversarial approaches such as TimeGAN Yoon et al. (2019) and QuantGAN Wiese et al. (2020) demonstrated that generative adversarial networks (GANs) can reproduce heavy tails, volatility clustering and cross-sectional dependence. Specialised variants soon followed: CorrGAN targets realistic correlation matrices for risk-parity and stress-testing tasks Marti (2020); Sig-Wasserstein GAN leverages the path-signature $W_{1}$ metric for conditional sampling that adapts to prevailing market states Liao et al. (2024); and Tail-GAN, trained with a VaR/ES-based objective that explicitly focuses learning on severe left-tail outcomes, enriches downside stress scenarios beyond what standard GANs provide Cont et al. (2022). These models supply high-resolution scenario sets that far outstrip traditional resampling in capturing dependence structures, regime shifts and tail behaviour, but they remain prone to training instability Mescheder et al. (2018) and offer no tractable likelihoods for rigorous evaluation.
@@
The state of the art has therefore shifted toward score-based diffusion models, which combine stable optimisation with likelihoods that are at least numerically tractable and come with explicit error bounds. Denoising Diffusion Probabilistic Models (DDPM) by Ho et al. (2020b), together with the continuous-time score-SDE formulation of Song et al. (2020), supply the theoretical foundation; in practice, likelihoods are obtained via discretised importance sampling rather than closed form, yet remain usable for model comparison. Finite-sample convergence guarantees—expressed as Wasserstein or KL error bounds under Lipschitz score error—have since been derived by De Bortoli (2022) and Gao et al. (2025), providing theoretical guarantees that are currently unavailable for standard GAN objectives.
@@
Building on these ideas, TimeGrad employs an autoregressive denoising architecture for multivariate prediction Rasul et al. (2021); ScoreGrad adapts continuous energy-based diffusion for improved sampling efficiency Yan et al. (2021); and CSDI applies conditional score matching to time series imputation Tashiro et al. (2021). Subsequent advances such as TimeDiff Shen and Kwok (2023) and TSDiff Kollovieh et al. (2023) show that diffusion generators can outperform transformer baselines in forecasting, while FTS-Diffusion Huang et al. (2024a) focuses on high-fidelity unconditional synthesis and demonstrates downstream forecasting gains through data augmentation rather than direct head-to-head benchmarks. Compared with GANs, diffusion models thus offer tractable log-likelihoods, stronger convergence guarantees, and greater robustness, making them particularly well suited to the dynamic mean–variance problems addressed in the present work.
@@
Generative models are moving beyond passive scenario engines and into the decision loop itself. Auxiliary-classifier GANs that condition on market states can learn return distributions whose samples feed directly into risk–return objectives, yielding portfolios that beat classical bootstrap or historical-window benchmarks Kim and Lee (2023). Diffusion-based methods advance the idea further: denoising-diffusion probabilistic models have been trained to generate realistic high-dimensional correlation matrices, enabling more faithful risk-parity and stress-testing exercises than traditional parametrics Kubiak et al. (2024).
@@
1.3 Notations
In this work, the data is in the form of time series with length $T\in\mathbb{N}$ and dimension $d\in\mathbb{N}$. We denote by $t\in\{1,2,\cdots,T\}$ the time index for the time-series. For general theoretical results, we use $X^{1:t}:=(X^{1},X^{2},\cdots,X^{t})$ ($Y^{1:t}:=(Y^{1},Y^{2},\cdots,Y^{t}$)) to denote random variables sampled from data distribution (approximated model, respectively), and $x^{1:t}$ ($y^{1:t}$) the realization. For portfolio selection problem, we use $S^{1:t}$ ($\tilde{S}^{1:t}$) and $s^{1:t}$ ($\tilde{s}^{1:t}$) instead. For diffusion model, we use $\tau\in[0,\mathcal{T}]$ to denote the diffusive time, i.e., the time flow of the forward and reversed SDE, with $\mathcal{T}>0$ the maximal diffusive time.
For a probability $\mathbb{P}$ on the space of time-series data $\mathbb{R}^{dT}$, $\mathbb{P}_{1:t}$ is the joint distribution of $(X^{1},X^{2},\cdots X^{t})$, thus $\mathbb{P}_{1:T}=\mathbb{P}$. Moreover, $\mathbb{P}_{x^{1:t}}$ is the conditional distribution of $X^{t+1}$, conditioning on $X^{1:t}=x^{1:t}$. By convention we use $\mathbb{P}$ to denote the original (real) model, hence the distribution of $X^{1:T}$ ($S^{1:T}$) and $\mathbb{Q}$ the alternative model, hence the distribution of $Y^{1:T}$ ($\tilde{S}^{1:T}$).
To present the quantification error bound of the score-based diffusion model, we use the adapted Wasserstein metric. More specifically, for two probability measures $\mathbb{P}$ and $\mathbb{Q}$ on $\mathbb{R}^{dT}$, a coupling $\pi$ between $\mathbb{P}$ and $\mathbb{Q}$ is said to be causal in $x$ if
$\pi(Y^{t+1}\in\mathrm{d}y^{t+1}|X^{1:T}=x^{1:T})=\pi(Y^{t+1}\in\mathrm{d}y^{t+1}|X^{1:t}=x^{1:t}).$ (1.4)
A coupling $\pi$ is said to be bicausal if it is causal in both $x$ and $y$ (c.f. Backhoff et al. (2017), Bayraktar and Han (2023)). Define $\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$ to be the set of all bicausal couplings between $\mathbb{P}$ and $\mathbb{Q}$. With the choice of transport cost $|x^{1:T}-y^{1:T}|:=(\sum_{t=1}^{T}|x^{t}-y^{t}|^{2})^{1/2}$, we define the adapted Wasserstein metric as follows:

### v3

@@ Abstract
In this paper, we tackle the dynamic mean-variance portfolio selection problem in a model-free manner, based on (generative) diffusion models. We propose using data sampled from the real model $\mathbb{P}$ (which is unknown) with limited size to train a generative model $\mathbb{Q}$ (from which we can easily and adequately sample). With adaptive training and sampling methods that are tailor-made for time-series data, we obtain quantification bounds between $\mathbb{P}$ and $\mathbb{Q}$ in terms of the adapted Wasserstein metric $\mathcal{A}\mathcal{W}_{2}$. Importantly, the proposed adapted sampling method also facilitates conditional sampling. In the second part of this paper, we provide the stability of the mean-variance portfolio optimization problems in $\mathcal{A}\mathcal{W}_{2}$. Then, combined with the error bounds and the stability result, we propose a policy gradient algorithm based on the generative environment, in which our innovative adapted sampling method provides approximate scenario generators. We illustrate the performance of our algorithm on both simulated and real data. For real data, the algorithm based on the generative environment produces portfolios that beat several important baselines, including the Markowitz portfolio, the equal weight (naive) portfolio, and S&P 500.
Subject Classification: dynamic programming: optimal control, finance: portfolio, computer science: artificial intelligence.
@@ 1 Introduction
Portfolio selection, a crucial and fundamental task in financial engineering, has been a subject of both theoretical and practical interest since the classical work Markowitz (1952). With the ability to account for intertemporal hedging and rebalancing in a non-stationary market, formulating portfolio selection as dynamic multi-period problems has become increasingly appealing; see Mossin (1968), Merton (1971) for early works. Researchers have explored this problem under various specifications, such as continuous-time dynamic mean-variance problems, where return models are prescribed by some stochastic differential equations Zhou and Li (2000), or discrete-time problems with independent intertemporal returns Li and Ng (2000). However, due to incompleteness of the information, estimation errors, or simply the infeasibility of fully modeling market uncertainty, we lack access to the real model, as opposed to model-based approaches. Instead, we are restricted to a limited set of financial data samples. Consequently, investigating portfolio selection problems in a model-free manner has been a long-standing topic in financial engineering; see discussions of ‘universal portfolio’ Cover (1991), stochastic portfolio theory Karatzas and Ruf (2017), Cuchiero et al. (2019), and the application of machine learning Ban et al. (2018), Gu et al. (2020). However, dynamic portfolio selection problems are rather involved, even with model-based approaches, due to their inherent intertemporal structure Campbell and Viceira (1999), high dimensionality Brandt et al. (2005), and time inconsistency Basak and Chabakauri (2010), Björk et al. (2014). Therefore, model-free dynamic portfolio selection problems are worth exploring.
@@
To address this challenge, we propose to leverage score-based diffusion models. These models can generate additional samples that exhibit distributional properties similar to the original data. Specifically, when the model $s^{1:T}\sim\mathbb{P}$ is unknown, we propose a data-driven approach to train a model $\mathbb{Q}$ that closely resembles $\mathbb{P}$. We utilize the model $\mathbb{Q}$ as an oracle in the subsequent policy gradient algorithm: we sample a large number of paths $\tilde{s}^{1:T}\sim\mathbb{Q}$ and train a reinforcement learning (RL) agent using these simulated paths. Theoretically, we show that our method achieves near-optimality in terms of Markowitz’s mean-variance criterion, thanks to two key contributions of the present paper: (1) a quantification bound between $\mathbb{P}$ and $\mathbb{Q}$, and (2) the model stability. From an engineering perspective, the approximated model $\mathbb{Q}$ consists of a score network $s_{\theta}$, which is the core of the score-matching technique used in generative models, and a Recurrent Neural Network (RNN) encoder $R_{\theta}$, which encodes the current state of the market. These two networks are trained simultaneously using historical price data. The reinforcement learning agent is trained on the generated data from $\mathbb{Q}$, which are produced by evaluating $s_{\theta}$ and $R_{\theta}$ in an adaptive manner. Finally, the agent, fed with the encoded (and updated) market state, outputs portfolio selection actions whose performance is evaluated through real-world experiments. For a summary of the theoretical results, see Figure 1, and for the workflow when implementing the proposed solution for model-free dynamic portfolio selection, refer to Figure 2.
[…]
@@ 1.1 Background, main results and contributions
The score matching technique, which has been a powerful tool in statistics since the fundamental work Hyvärinen (2005), has also been incorporated into generative models in various works, including Sohl-Dickstein et al. (2015), Song and Ermon (2019), and others. Specifically, Song et al. (2021b) unifies the concept of score-based generative models and diffusion probabilistic models (Ho et al. (2020)) into a single framework, known as score-based diffusion models, which we adopt in this work.
@@
where $p(\tau,\cdot)$ denotes the density of $X_{\tau}$. Because $p_{\rm data}$ is unknown, so is $p(\tau,\cdot)$. Under the score‑matching theory, an accurate estimate of the score $\nabla\log p$ guarantees that, for sufficiently large $\mathcal{T}$, the law of $Y_{\mathcal{T}}$ approximates $p_{\rm data}$ under various metrics; see Tang and Zhao (2025) for a comprehensive survey and Gao et al. (2025) for recent Wasserstein bounds. Score-based diffusion models have found numerous applications across a wide range of areas, including image and video generation (Rombach et al. (2022), Melnik et al. (2024)), molecule simulation Hoogeboom et al. (2022) and health Abramson et al. (2024). A very recent work Chen et al. (2025) has also specifically applied the score-based diffusion model to financial data generation, exploiting low-dimensional structures in high-dimensional return data.
@@
Nevertheless, because we are interested in dynamic decision-making problems, the data take the form of time-series: $X=(X^{1},X^{2},\cdots,X^{T})=X^{1:T}$. For instance, $T=12$ and each $X^{t}$ is the monthly price data, then $X$ is a one-year price path. Although score-based diffusion models have been applied to time-series prediction (see Lin et al. (2024) for a survey), to the best of our knowledge, there is still no theoretical result that establishes foundation for time-series data generation. Treating the state space naively as the enlarged Euclidean space $\mathbb{R}^{dT}$ is unsatisfactory: it overlooks the intertemporal structure and the non-anticipation requirements inherent in dynamic problems. Indeed, existing results show that, for dynamic settings, value functions are unstable under standard probabilistic metrics such as Wasserstein; see, e.g., Backhoff-Veraguas et al. (2020). This observation motivates the use of the adapted Wasserstein metric, $\mathcal{A}\mathcal{W}_{2}$, which is better suited to time-series data and dynamic tasks; see Backhoff-Veraguas et al. (2020) for definitions. Furthermore, a dynamic control problem requires sampling not only from the joint distribution of $\mathbb{P}$ but also from the conditional distribution $\mathbb{P}(\cdot|X^{1:t}=x^{1:t})$ given observations up to time $t$. Such conditional sampling is indispensable for our subsequent policy-gradient algorithm (i.e., for training the RL agent). Traditional score-based diffusion models, however, do not support this functionality. Consequently, we must develop new training and sampling procedures that (i) admit theoretical score-matching error bounds under the adapted Wasserstein metric and (ii) enable conditional sampling for any observation input.
@@
Our first set of results meets these two requirements. For notational convenience, we fix $\beta\equiv 2$ in (1.1) and, for each $t=1,2,\dots,T-1$, consider the following conditional VP-SDE:
$\begin{cases}&\mathrm{d}X^{t+1}_{\tau}=-X^{t+1}_{\tau}\mathrm{d}\tau+\sqrt{2}\mathrm{d}B^{t+1}_{\tau},\\ &X^{t+1}_{0}\sim\mathbb{P}_{x^{1:t}},\end{cases}$ (1.3)
@@
where $\mathbb{P}_{x^{1:t}}$ denotes the law of $X^{t+1}$ conditional on $X^{1:t}=x^{1:t}$. This formulation allows us to define the conditional score function $s^{t+1}(\tau,x^{1:t},x)=\nabla_{x}\log p_{t+1}(\tau,x^{1:t},x)$, where $p_{t+1}(\tau,x^{1:t},\cdot)$ is the density of $X^{t+1}_{\tau}$ in (1.3). Assume that we have an approximator $\hat{s}^{t+1}$ that is close to $s$ in a suitable $L^{2}$ sense. We generate a new trajectory $y^{1:T}$ adaptively: to obtain $y^{1}$ we start from pure noise and run the reverse VP-SDE (1.2), setting $y^{1}=Y_{\mathcal{T}}$; to obtain $y^{t+1}$ we use the previously generated path $y^{1:t}$ as conditional input and run (1.2) with the score replaced by $\hat{s}^{t+1}(\cdot,y^{1:t},\cdot)$. See Algorithm 1 for more details. Under appropriate technical conditions, we obtain the following bound:
@@
Denote by $\mathbb{Q}^{\mathcal{T}}$ the distribution of $y^{1:T}$. Then $\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq E(\mathcal{T},\varepsilon_{\rm score})$, where $\varepsilon_{\rm score}$ is the score-matching error between $\hat{s}$ and $s$. Moreover, $E(\mathcal{T},\varepsilon_{\rm score})$ has an explicit form and can be made arbitrarily small by taking $\mathcal{T}$ sufficiently large and $\varepsilon_{\rm score}$ sufficiently small.
@@
While we can now generate new samples from $\mathbb{Q}$, several steps remain before we can fully address the dynamic portfolio selection problem. First, because we now work under the surrogate model $\mathbb{Q}$, we must establish model stability: even if $\mathbb{P}$ and $\mathbb{Q}$ are close and we can solve the mean–variance problem under $\mathbb{Q}$, does the optimizer obtained from $\mathbb{Q}$ remain near‑optimal for the original model $\mathbb{P}$? We answer this question in the affirmative:
@@
    We extend the score‑based diffusion‑model framework to time‑series data. Using a novel adaptive training and sampling scheme, we establish theoretical error bounds under the adapted Wasserstein metric. The framework also supports conditional sampling, which is crucial for solving model‑free portfolio‑selection problems and constructing high‑performance portfolio strategies;
@@ 1.2 Related works
The use of simulated price paths to inform investment and corporate‐finance decisions predates modern portfolio theory. Hertz (1964) first popularized Monte Carlo simulation for capital‐budgeting analysis, and Boyle (1977) introduced path‐wise Monte Carlo methods to option valuation. During the 1970s–1990s, more sophisticated parametric dynamics, jumping diffusion Merton (1976), stochastic volatility Heston (1993), and SV with jumps Bates (1996) - made large-scale pathwise risk analysis computationally tractable.
@@
As asset universes became high-dimensional and tail dependencies more intricate, researchers began exploring fully data-driven generators. Early adversarial approaches such as TimeGAN Yoon et al. (2019) and QuantGAN Wiese et al. (2020) demonstrated that generative adversarial networks (GANs) can reproduce heavy tails, volatility clustering and cross-sectional dependence. Specialised variants soon followed: CorrGAN targets realistic correlation matrices for risk-parity and stress-testing tasks Marti (2020); Sig-Wasserstein GAN leverages the path-signature $W_{1}$ metric for conditional sampling that adapts to prevailing market states Liao et al. (2024); VolGAN specializes in generating implied volatility surfaces and is applied in constructing hedging portfolio Vuletić and Cont (2024), Cont and Vuletić (2025); and Tail-GAN, trained with a VaR/ES-based objective that explicitly focuses learning on severe left-tail outcomes, enriches downside stress scenarios beyond what standard GANs provide Cont et al. (2025). These models supply high-resolution scenario sets that far outstrip traditional resampling in capturing dependence structures, regime shifts and tail behaviour, but they remain prone to training instability Mescheder et al. (2018) and offer no tractable likelihoods for rigorous evaluation.
@@
The state of the art has therefore shifted toward score-based diffusion models, which combine stable optimisation with likelihoods that are at least numerically tractable and come with explicit error bounds. Denoising Diffusion Probabilistic Models (DDPM) by Ho et al. (2020), together with the continuous-time score-SDE formulation of Song et al. (2021b), supply the theoretical foundation; in practice, likelihoods are obtained via discretised importance sampling rather than closed form, yet remain usable for model comparison. Finite-sample convergence guarantees—expressed as Wasserstein or KL error bounds under Lipschitz score error—have since been derived by Bortoli (2022) and Gao et al. (2025), providing theoretical guarantees that are currently unavailable for standard GAN objectives.
@@
Building on these ideas, TimeGrad employs an autoregressive denoising architecture for multivariate prediction Rasul et al. (2021); ScoreGrad adapts continuous energy-based diffusion for improved sampling efficiency Yan et al. (2021); and CSDI applies conditional score matching to time-series imputation Tashiro et al. (2021). Subsequent advances such as TimeDiff Shen and Kwok (2023) and TSDiff Kollovieh et al. (2023) show that diffusion generators can outperform transformer baselines in forecasting, while FTS-Diffusion Huang et al. (2024a) focuses on high-fidelity unconditional synthesis and demonstrates downstream forecasting gains through data augmentation rather than direct head-to-head benchmarks. Compared with GANs, diffusion models thus offer tractable log-likelihoods, stronger convergence guarantees, and greater robustness, making them particularly well suited to the dynamic mean–variance problems addressed in the present work.
@@
Generative models are moving beyond passive scenario engines and into the decision loop itself. Auxiliary-classifier GANs that condition on market states can learn return distributions whose samples feed directly into risk–return objectives, yielding portfolios that beat classical bootstrap or historical-window benchmarks Kim and Lee (2023). Diffusion-based methods advance the idea further: denoising-diffusion probabilistic models have been trained to generate realistic high-dimensional correlation matrices, enabling more faithful risk-parity and stress-testing exercises than traditional parametrics Kubiak et al. (2024), and score-based (continuous-time) diffusion model admits explicit statistical guarantee that addressing curse of dimensionality when there are hidden low-dimensional structure Chen et al. (2025).
@@
1.3 Organization of the paper
The rest of this paper is organized as follows. In Section 2 we present an adaptive training and sampling framework of score-based diffusion models, and provide a $\mathcal{A}\mathcal{W}_{2}$-quantification error bound between the output distribution and data distribution. In Section 3, we turn to a specific mean-variance portfolio selection problem, and establish the model stability in terms of the adapted Wasserstein metric (i.e., the value function is Lipschitz in the adapted Wasserstein metric). Section 4 focuses on applications of the theoretical results and, in particular, develops a policy gradient algorithm based on generative environment along with experimental results. 5 concludes the main context. Appendices include technical discussions and proofs of the main results in Section 2.
We will close this section by introducing frequently used notation in the paper.
1.4 Frequently used notation
In this work, the data is in the form of time-series with length $T\in\mathbb{N}$ and dimension $d\in\mathbb{N}$. We denote by $t\in\{1,2,\cdots,T\}$ the time index for the time-series. For general theoretical results, we use $X^{1:t}:=(X^{1},X^{2},\cdots,X^{t})$ ($Y^{1:t}:=(Y^{1},Y^{2},\cdots,Y^{t}$)) to denote random variables sampled from data distribution (approximated model, respectively), and $x^{1:t}$ ($y^{1:t}$) the realization. For portfolio selection problem, we use $S^{1:t}$ ($\tilde{S}^{1:t}$) and $s^{1:t}$ ($\tilde{s}^{1:t}$) instead. For the diffusion model, we use $\tau\in[0,\mathcal{T}]$ to denote the diffusive time, i.e., the time flow of the forward and reversed SDE, with $\mathcal{T}>0$ the maximal diffusive time.
For a probability $\mathbb{P}$ on the space of time-series data $\mathbb{R}^{dT}$, $\mathbb{P}_{1:t}$ is the joint distribution of $(X^{1},X^{2},\cdots X^{t})$, thus $\mathbb{P}_{1:T}=\mathbb{P}$. Moreover, $\mathbb{P}_{x^{1:t}}$ is the conditional distribution of $X^{t+1}$, conditioning on $X^{1:t}=x^{1:t}$. By convention, we use $\mathbb{P}$ to denote the original (real) model, hence the distribution of $X^{1:T}$ ($S^{1:T}$) and $\mathbb{Q}$ the alternative model, hence the distribution of $Y^{1:T}$ ($\tilde{S}^{1:T}$).
To present the quantification error bound of the score-based diffusion model, we use the adapted Wasserstein metric. More specifically, for two probability measures $\mathbb{P}$ and $\mathbb{Q}$ on $\mathbb{R}^{dT}$, a coupling $\pi$ between $\mathbb{P}$ and $\mathbb{Q}$ is said to be causal in $x$ if for any $t\in\{1,2,\cdots,T\}$,
$\pi(Y^{t}\in\mathrm{d}y^{t}|X^{1:T}=x^{1:T})=\pi(Y^{t}\in\mathrm{d}y^{t}|X^{1:t}=x^{1:t}).$ (1.4)
A coupling $\pi$ is said to be bicausal if it is causal in both $x$ and $y$ (c.f. Backhoff et al. (2017), Bayraktar and Han (2025+)). Define $\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$ to be the set of all bicausal couplings between $\mathbb{P}$ and $\mathbb{Q}$. With the choice of transport cost $|x^{1:T}-y^{1:T}|:=(\sum_{t=1}^{T}|x^{t}-y^{t}|^{2})^{1/2}$, we define the adapted Wasserstein metric as follows: